If 3 < x < 7 and 4 < y < 7, which of the following best describes the range of values of x – y?
A. –4 < x – y < 3
B. 0 < x – y < 4
C. 3 < x – y < 4
D. 3 < x – y < 7
E. 4 < x – y < 7

Answer by richwmiller(17219)

Try some numbers
Let's take the extremes of x and y
x=4 y=6 x-y=-2
x=6 y=4 x-y=2
x=4 y=5 x-y=-1
x=6 y=6 x-y=0
So we see that x-y can be less than zero
so A) is the only one that matches